for both the diode D1 and D2 the forward resistance is 10 ohm and cut in voltage is 0.5V. Calculate the current through 20 ohm resistance connected in the circuit.
1
Expert's answer
2020-04-15T01:51:44-0400

Cut-in voltage or knee voltage or threshold voltage is the minimum voltage( in forward bias condition)after which the diode current rises rapidly and diode is said to be conducting

And here we are given that the diode forward biased resistance is 10 ohm and the external resistance is 20 ohm

current = "\frac{V-V_{cut\ in}}{R_{total}}=\frac{V-0.5}{10+20}=\frac{V-0.5}{30}"